Use appropriate algebra and Theorem 7.2.1 to fi S L-1 {3+ s2 + 35 - 4 х Need Help? Read it Talk to a Tutor
Answer #1

Given F(s)=\frac{s}{s^2+3s-4}

\frac{s}{s^2+3s-4}=\frac{s}{(s-1)(s+4)}

By using partial fractions, divide the above function as:

\frac{s}{(s-1)(s+4)}=\frac{A}{s-1}+\frac{B}{s+4}

\Rightarrow s=A(s+4)+B(s-1)

By equating the coefficients, we get

A+B=1,

4A-B=0

By solving these 2 equations, we get:

A=1/5, B=4/5

\frac{A}{s-1}+\frac{B}{s+4}=\frac{1/5}{s-1}+\frac{4/5}{s+4}

Now finding inverse Laplace transform of F(s):

L^{-1}\left ( \frac{s}{s^2+3s-4}\right )=L^{-1}\left \{ \frac{1/5}{s-1} \right \}+L^{-1}\left \{ \frac{4/5}{s+4} \right \}. =\frac{1}{5}L^{-1}\left \{ \frac{1}{s-1} \right \}+\frac{4}{5}L^{-1}\left \{ \frac{1}{s+4} \right \}

=\frac{1}{5}e^t+\frac{4}{5}e^{-4t}